# Markov switching zero-inflated space-time multinomial models for comparing multiple infectious diseases

### v1

@@ Abstract
Despite multivariate spatio-temporal counts often containing many zeroes, zero-inflated multinomial models for space-time data have not been considered. We are interested in comparing the transmission dynamics of several co-circulating infectious diseases across space and time where some can be absent for long periods. We first assume there is a baseline disease that is well-established and always present in the region. The other diseases switch between periods of presence and absence in each area through a series of coupled Markov chains, which account for long periods of disease absence, disease interactions and disease spread from neighboring areas. Since we are mainly interested in comparing the diseases, we assume the cases of the present diseases in an area jointly follow an autoregressive multinomial model. We use the multinomial model to investigate whether there are associations between certain factors, such as temperature, and differences in the transmission intensity of the diseases. Inference is performed using efficient Bayesian Markov chain Monte Carlo methods based on jointly sampling all presence indicators. We apply the model to spatio-temporal counts of dengue, Zika and chikungunya cases in Rio de Janeiro, during the first triple epidemic.
@@ 1 Introduction
Comparing the transmission dynamics of several co-circulating infectious diseases is an important problem in epidemiology. In this paper, we are specifically concerned with comparing the transmission dynamics of the vector-borne diseases Zika, chikungunya and dengue. Since 2015 several countries in Latin America have experienced triple epidemics of these diseases, placing an immense burden on the local populations (Rodriguez-Morales et al., 2016; Bisanzio et al., 2018). Despite being transmitted by the same vector, the Aedes aegypti, several studies have found spatio-temporal differences in clusters of the three diseases (Freitas et al., 2019; Kazazian et al., 2020). Investigating how these differences may be related to covariates could help more efficiently distribute medical resources, as health interventions can vary between the diseases (Freitas et al., 2019). Schmidt et al. (2022) attempted this using a multinomial model, but they only investigated how the relative distribution of the diseases varied spatially. There is also an interest in seeing how the transmission of the three diseases may differ by spatio-temporal factors. For instance, laboratory studies have shown that the transmission of Zika may be more sensitive to temperature compared to dengue, which is important as it implies Zika will not be able to spread as effectively into North America and the highlands (Tesla et al., 2018).
As a motivating example, we analyze bi-weekly reported case counts of Zika, chikungunya, and dengue in the 160 neighborhoods of Rio de Janeiro, Brazil, between 2015-2016, during the first triple epidemic there. Since we are mainly interested in comparing the diseases, we condition on the total number of cases in each area and bi-week and use a multinomial distribution to model their relative allocations (Dabney and Wakefield, 2005; Dreassi, 2007; Schmidt et al., 2022). Several authors have proposed modeling multinomial data across time, space, or both by adding correlated random effects to multinomial link functions, such as the log relative odds (Pereira et al., 2018; Sosa et al., 2023). Proposed distributions for the random effects include conditional autoregressive distributions (Schmidt et al., 2022), multivariate logit-beta distributions (Bradley et al., 2019) and multivariate normal random walks (Cargnoni et al., 1997). Autoregressive approaches (Loredo-Osti and Sutradhar, 2012) have also been popular, where the current multinomial probabilities in an area are assumed to depend on past observations in the area (Fokianos and Kedem, 2003) and neighboring areas (Tepe and Guldmann, 2020). Since new infectious disease cases are transmitted from previous cases, conditioning on past observations is often a natural way to model infectious disease counts across time and space (Bauer and Wakefield, 2018). Therefore, we primarily take an autoregressive-driven approach in this paper. We assume that if conditions are more favorable for one disease over the others, then the share of that disease will grow over time. We represent the relative favorability of the diseases through a series of autoregressive parameters that are linked to covariates. This allows us to investigate whether a covariate is associated with conditions that are more favorable toward the transmission of one of the diseases. As we will show, our approach is epidemiologically motivated and can be derived from a series of Reed-Frost chain binomial susceptible-infectious-recovered (SIR) models (Abbey, 1952; Vynnycky, 2010; Bauer and Wakefield, 2018).
@@
A potential issue with applying a multinomial model to multivariate spatio-temporal infectious disease counts is that many infectious diseases, especially vector-borne diseases, can be absent for long periods of time in an area (Bartlett, 1957; Coutinho et al., 2006; Adams and Boots, 2010). This behavior is illustrated in Figure 1 which shows bi-weekly reported case counts of dengue, Zika and chikungunya in the western Rio neighborhood of Praça Seca, between 2015-2016. While dengue cases are reported consistently throughout the two years, very few Zika cases and no chikungunya cases are reported in the first year. These patterns are typical of most neighborhoods. If a disease is absent in an area, then there should be no chance of cases being reported. However, a multinomial model will always assign a positive probability to cases being reported, which may lead to many more zeroes in the counts than can be realistically produced by the fitted model (Koslovsky, 2023). Also, we do not want to compare the transmission dynamics of a present disease and an absent one as it would likely overestimate the favorability of the present disease.
@@
To account for long periods of disease absence, we will place our proposed autoregressive multinomial model within a zero-inflated multinomial modeling framework (Zeng et al., 2022; Koslovsky, 2023). Zero-inflated multinomial models allow categories, in our case diseases, to be absent from observations. If a category is absent, its multinomial probability is fixed at 0. The presence of a category is usually treated as random and generated through an independent Bernoulli process (Koslovsky, 2023). While a zero-inflated multinomial model would allow diseases to be absent in areas, existing approaches have made two independence assumptions that are inappropriate for spatio-temporal infectious disease counts. Firstly, they have assumed that the multivariate counts are independent conditional on the present categories of each observation. Infectious disease counts usually exhibit a high amount of correlation across space and time due to disease transmission (Bauer and Wakefield, 2018). Secondly, past approaches assumed that if a category is present in one observation it does not affect whether that category or other categories are present in other observations (independence of categorical presence). If an infectious disease is present in an area, then it is more likely to be present in the next time period, as disease presence and absence are usually persistent (Coutinho et al., 2006). Also, the presence of one disease may affect whether other diseases are present due to potential disease interactions (Paul et al., 2008). Finally, due to between-area disease spread (Grenfell et al., 2001), the status of the disease in an area will affect its presence in neighboring areas.
@@
To address the above issues, we first choose a baseline disease that is well-established in the region and assume it is always present. This avoids all diseases being absent in an area at once which would be impossible if the total number of cases were not 0. We assume all other diseases switch back and forth between periods of presence and absence in each area through a series of Markov chains. The Markov chains account for long periods of disease absence, disease interactions and covariates, including the prevalence of the diseases in neighboring areas. For the diseases present in an area together, we assume that the relative distribution of their cases follows our proposed space-time autoregressive multinomial model. The absent diseases always report 0 cases. The rest of this paper is organized as follows. Section 2 introduces our proposed autoregressive multinomial model for comparing the transmission dynamics of multiple infectious diseases in a space-time setting. Section 2.1 extends the model to incorporate zero inflation and deal with long periods of disease absence. Section 3 details our Bayesian inferential procedure which utilizes the forward filtering backward sampling (FFBS) algorithm of Chib (1996) for efficient inference. In Section 4, we apply our model to cases of dengue, Zika and chikungunya in Rio de Janeiro during the first triple epidemic there. We close with a general discussion in Section 5.
@@ 2 An Autoregressive Multinomial Model for Comparing Transmission Dynamics
Assume we have reported case counts for $K$ different infectious diseases across $i=1,\dots,N$ areas and $t=1,\dots,T$ time periods. Let $\bm{y}_{it}=(y_{1it},\dots,y_{Kit})^{T}$ represent the vector of case counts for all diseases in area $i$ during the time interval $(t-1,t]$, where $y_{kit}$ represents the reported case counts for disease $k$. We will first present our model without zero inflation and then extend to the zero-inflated case in Section 2.1. Since we are mainly interested in comparing the diseases, we condition on the total number of cases and model $\bm{y}_{it}$ using a multinomial distribution,
@@
where $\text{total}_{it}=\sum_{k=1}^{K}y_{kit}$ and $\bm{y}_{t-1}=(\bm{y}_{1(t-1)},\dots,\bm{y}_{N(t-1)})^{T}$ is the vector of all case counts reported in the previous time, giving the model an autoregressive structure. In (1), $\bm{\pi}_{it}=(\pi_{1it},\dots,\pi_{Kit})^{T}$ represents the expected relative distribution of the disease counts at time $t$ in area $i$ given the previous observed counts $\bm{y}_{t-1}$. If conditions in $(t-1,t]$ favor some diseases over others then the share of the favored diseases should grow relative to those less favored. Therefore, we model the relative odds as,
@@
for $k=2,\dots,K$, where we assume disease 1 is the baseline disease and we add 1 to the top and bottom to avoid dividing by zero, adding such constants is common in autoregressive count models (Liboschik et al., 2017; Fritz and Kauermann, 2022). The parameters $\zeta_{k}$ and $\zeta_{1}$, assumed to be between 0 and 1, are meant to account for nonhomogeneous mixing (Wakefield et al., 2019). That is, as cases increase there tends to be a dampening effect on transmission as individuals avoid infection and governments implement interventions. In (2), $\lambda_{kit}>0$, which we link to covariates below, is a parameter that represents the favorability of conditions for the transmission of disease $k$ relative to the baseline disease during $(t-1,t]$. For instance, if $\lambda_{kit}>1$ ($\lambda_{kit}<1$) then we would expect the share of disease $k$ to grow (shrink) relative to the baseline disease in $(t-1,t]$, after adjusting for nonhomogeneous mixing with $\zeta_{k}$ and $\zeta_{1}$. If $\zeta_{k}\approx\zeta_{1}<<1$, like for our motivating example in Section 4, then this means that if $\lambda_{kit}>1$ the share of disease $k$ will grow initially from an equal share. However, as disease $k$ becomes more dominant the share will be pulled back towards equality as individuals avoid infection from the dominant disease.
@@
$\log(\lambda_{kit})=\alpha_{0ki}+\bm{x}_{kit}^{T}\bm{\alpha}_{k}+\phi_{kit},$ (3)
where $\alpha_{0ki}\sim N(\alpha_{0k},\sigma_{k}^{2})$ is a normal random intercept meant to account for between area differences and $\bm{x}_{kit}$ is a vector of covariates that might affect the favorability of disease $k$ relative to the baseline disease. We model the random effects $\phi_{kit}$ using a multivariate normal distribution (Xia et al., 2013; Zeng et al., 2022),
@@
where $\bm{\Sigma}$ is a $K-1$ by $K-1$ variance-covariance matrix. These random effects are mainly meant to account for overdispersion, which is a very common issue when modeling multinomial counts (Xia et al., 2013). They also allow for a more flexible correlation structure between the individual disease counts, see the supplementary materials (SM) Section 2.
@@
There is an epidemiological interpretation of the parameters in Equations (1)-(4). As we show in SM Section 1, if all diseases follow a Reed-Frost chain binomial SIR model (Abbey, 1952; Vynnycky, 2010; Bauer and Wakefield, 2018) with roughly the same serial interval, the time between successive generations of cases, then we can derive Equations (1)-(4). Under this derivation, $\lambda_{kit}$ in (2) represents the ratio of the effective reproduction number of disease $k$ and the baseline disease. The effective reproduction number is an important measure of transmission in epidemiology that gives the average number of new infections produced by a single infectious individual in the current population before they recover (Vynnycky, 2010). The covariate effect $\alpha_{kj}$ then represents the difference between the effect of covariate $x_{kitj}$ on the effective reproduction number of disease k and the baseline disease, see SM Section 1 for more details.
@@
While the Reed-Frost model is commonly used (Wakefield et al., 2019), it makes assumptions that are not appropriate for many diseases (Abbey, 1952), and the model can be sensitive to underreporting and reporting delay when it comes to estimating the effective reproduction number (Bracher and Held, 2021; Quick et al., 2021). Regardless, the Reed-Frost derivation does reveal an important source of potential confounding due to differences in the susceptible populations, i.e., the population of individuals who are not immune to the disease. From the Reed-Frost derivation, see SM Equation (8), ideally, we would add $\log(\delta_{ki(t-1)}/\delta_{1i(t-1)})$ to $\log(\lambda_{kit})$ in Equation (3), where $\delta_{ki(t-1)}$ is the size of the susceptible population for disease $k$ in area $i$ at time $t-1$. This makes sense intuitively, if a disease has a larger susceptible population compared to another then transmission for that disease will naturally be favored. The effect of any covariates correlated with the ratio of the susceptible populations could then be confounded. We conduct a sensitivity analysis to attempt to control for this for our motivating example in SM Section 8.
Finally, we note that the number of cases in neighboring areas may affect the favorability of the diseases as individuals will mix with those in neighboring areas (Grenfell et al., 2001). Therefore, we allow functions of disease cases in neighboring areas, such as the prevalence across neighboring areas, to be added to the covariate vector $\bm{x}_{kit}$. Also, we allow previous counts of other diseases, e.g., $\log(y_{ji(t-1)}+1)$ for $j\neq k$, to be added to $\bm{x}_{kit}$ to account for potential disease interactions (Freitas et al., 2019), see Section 4 for an example. We will call the model defined by (1)-(4) the autoregressive multinomial (ARMN) model.

### v2

@@ Abstract
Univariate zero-inflated models are increasingly being used to account for excess zeros in spatio-temporal infectious disease counts. However, the multivariate case is challenging due to the need to account for correlations across space, time and disease in both the count and zero-inflated components of the model. We are interested in comparing the transmission dynamics of several co-circulating infectious diseases across space and time, where some of the diseases can be absent for long periods. We first assume there is a baseline disease that is well-established and always present in the region. The other diseases switch between periods of presence and absence in each area through a series of coupled Markov chains, which account for long periods of disease absence, disease interactions and disease spread from neighboring areas. Since we are mainly interested in comparing the diseases, we assume the cases of the present diseases in an area jointly follow an autoregressive multinomial model. We use the multinomial model to investigate whether there are associations between certain factors, such as temperature, and differences in the transmission intensity of the diseases. Inference is performed using efficient Bayesian Markov chain Monte Carlo methods based on jointly sampling all unknown presence indicators. We apply the model to spatio-temporal counts of dengue, Zika, and chikungunya cases in Rio de Janeiro, during the first triple epidemic there.
@@ 1 Introduction
In epidemiology, there has been a growing interest in analyzing multivariate counts of co-circulating infectious diseases across space and time to compare transmission dynamics (Freitas et al., 2019). In this paper, we analyze bi-weekly reported case counts of Zika, chikungunya, and dengue in the 160 neighborhoods of Rio de Janeiro, Brazil, between 2015-2016, during the first triple epidemic there. The substantive aim of our analysis is to investigate how the transmission of the three diseases differs by factors such as temperature. For instance, if Zika transmits less effectively at lower temperatures compared to dengue, it would imply that its range of spread is more limited, which is helpful information for controlling the diseases (Tesla et al., 2018). Since we are mainly interested in comparing the diseases, we condition on the total number of cases in each area and bi-week and use a multinomial distribution to model their relative allocations (Dabney and Wakefield, 2005; Dreassi, 2007; Schmidt et al., 2022).
@@
A potential issue with applying a multinomial model to multivariate spatio-temporal infectious disease counts is that many infectious diseases, especially vector-borne diseases, can be absent for long periods of time in an area (Bartlett, 1957; Coutinho et al., 2006; Adams and Boots, 2010). This behavior is illustrated in Figure 1, which shows bi-weekly reported case counts of dengue, Zika, and chikungunya in the western Rio neighborhood of Praça Seca, between 2015-2016. While dengue cases are reported consistently throughout the two years, very few Zika cases and no chikungunya cases are reported in the first year. These patterns are typical of most neighborhoods. If a disease is absent in an area, then there should be no chance of cases being reported. However, a multinomial model will always assign a positive probability to cases being reported. Therefore, there will likely be many more zeros in the counts than can be realistically produced by a fitted multinomial model (we investigate this in more detail in Section 4) (Koslovsky, 2023). Generally, when the fitted model cannot produce enough zeros, we say that the data contains excess zeros, which is very common when modeling infectious disease counts across space and time (Arab, 2015). Also, we do not want to compare the transmission dynamics of a present disease and an absent one, as it would likely overestimate the favorability of the present disease.
@@
When analyzing a single disease, zero-inflated count models (Lambert, 1992) are often used to deal with excess zeros in spatio-temporal infectious disease counts (Fernandes et al., 2009; Arab, 2015; Wangdi et al., 2018). With these approaches, the presence/absence of the disease is first generated through a Bernoulli process (Fernandes et al., 2009). Then, if the disease is absent in an area, it is assumed that no cases will be reported. Conversely, if the disease is present, the number of reported cases is generated through a count process, typically negative binomial or Poisson (Aktekin and Musal, 2015). A zero that arises from the count process is often interpreted as the result of the disease going undetected, referred to as a false zero (Vergne et al., 2014). This is opposed to a true zero from the Bernoulli process, which corresponds to the actual absence of the disease. Random effects can be added to both the count and Bernoulli (zero-inflated) parts of the model to account for correlations across space and time (Torabi, 2017; Giorgi et al., 2018). For example, Hoef and Jansen (2007) included random walks to account for correlations across time and conditional autoregressive effects (Besag et al., 1991) to account for correlations across space in both parts.
While the above models are popular in the univariate case, very little attention has been given to multivariate spatio-temporal zero-inflated disease modeling in general. To the best of our knowledge, no zero-inflated multinomial models have been proposed for space-time data. Pavani and Moraga (2022) and Rotejanaprasert et al. (2021) both considered zero-inflated spatio-temporal multivariate Poisson models. They used spatial and temporal random effects to account for correlations in the data. There is a close relationship between multivariate Poisson and multinomial models for multivariate disease counts (Schmidt et al., 2022). However, we prefer the multinomial approach, as it eliminates many nuisance parameters if the interest is only in comparing the diseases; see the supplementary materials (SM) Section 1. In addition, the above multivariate Poisson proposals have some important limitations that make them unsuitable for our application, including not allowing the probability of disease presence to change over time. Vector-borne diseases are highly seasonally driven and are more likely to be present in the summer and absent in the winter (Coutinho et al., 2006). Also, if an infectious disease is present in an area, it should be more likely to be present in that area and neighboring areas in the next time (Douwes-Schultz and Schmidt, 2022). The challenge with scaling up the univariate random effect models described in the previous paragraph to the multivariate case is the need to include random effects correlated across space, time and disease (not just space and time) in both the count and zero-inflated parts of the model, which would be computationally challenging.
@@
To account for long periods of disease absence, we will extend the zero-inflated multinomial modeling framework of Zeng et al. (2022) to accommodate spatio-temporal infectious disease data. Similar to the univariate models described above, zero-inflated multinomial models allow categories (in our case diseases) to be absent from observations. If a category is absent, its multinomial probability is fixed at 0, meaning the count for that category will be 0. The presence of the categories is usually treated as random and generated through a series of independent Bernoulli processes (Koslovsky, 2023). While a zero-inflated multinomial model would allow some diseases to be absent in certain areas, existing approaches have made two independence assumptions that are inappropriate for spatio-temporal infectious disease counts. Firstly, they have assumed that the multivariate counts are independent conditional on the present categories of each observation. Infectious disease counts usually exhibit a high amount of correlation across space and time due to disease transmission (Bauer and Wakefield, 2018). Secondly, past approaches assumed that if a category is present in one observation, it does not affect whether that category or other categories are present in other observations (independence of categorical presence). If an infectious disease is present in an area, then it is more likely to be present in the next time period, as disease presence and absence are usually persistent (Coutinho et al., 2006). Also, the presence of one disease may affect whether other diseases are present due to potential disease interactions (Paul et al., 2008). Finally, due to between-area disease spread (Grenfell et al., 2001), the status of the disease in an area will affect its presence in neighboring areas.
To address the above issues, we first choose a baseline disease that is well-established in the region and assume it is always present. This avoids all diseases being absent in an area at once, which would be impossible if the total number of cases were not 0. As we discuss in Section 2.1, this problem was not considered by Zeng et al. (2022). We then assume that all other diseases switch back and forth between periods of presence and absence in each area through a series of nonhomogeneous coupled Markov chains. That is, we allow the probabilities of disease presence to depend on whether the disease was present previously in the area; whether other diseases were present previously, to account for interactions; covariates, like temperature; and the prevalence of the disease in neighboring areas, to account for spread. Therefore, unlike Pavani and Moraga (2022) and Rotejanaprasert et al. (2021), we allow the risk of disease presence to change dynamically across space, time and disease. Also, unlike Zeng et al. (2022), the presence of the categories is not independent between observations. For instance, if a disease is present in an area, this disease, and potentially other diseases, can be more likely to be present again. For the count part of our zero-inflated model, we assume that if a disease is absent in an area, no cases will be reported. If a disease is present, we assume that the odds of a new case being from that disease relative to the baseline disease (relative odds) follows an autoregressive log-linear model. We allow the log relative odds to depend on past counts in the area and neighboring areas (Fokianos and Kedem, 2003; Tepe and Guldmann, 2020). Therefore, unlike Zeng et al. (2022), we do not assume the multivariate counts are independent conditional on the presence of the diseases, and we account for the transmission of the present diseases between neighboring areas (Bauer and Wakefield, 2018). Finally, our model can be written as an autoregressive Markov switching model, meaning Bayesian inference is very tractable through the forward filtering backward sampling algorithm (Frühwirth-Schnatter, 2006).
@@
The rest of this paper is organized as follows. Section 2 introduces our proposed autoregressive multinomial model for comparing the transmission dynamics of multiple infectious diseases in a space-time setting. Section 2.1 extends the model to incorporate zero inflation and deal with long periods of disease absence. Section 3 details our Bayesian inferential procedure, which utilizes the forward filtering backward sampling (FFBS) algorithm of Chib (1996) for efficient inference. In Section 4, we apply our model to cases of dengue, Zika and chikungunya in Rio de Janeiro during the first triple epidemic there. We close with a general discussion in Section 5.
@@ 2 An Autoregressive Multinomial Model for Comparing Transmission Dynamics
Assume we have reported case counts for $K$ different infectious diseases across $i=1,\dots,N$ areas and $t=1,\dots,T$ time periods. Let $\bm{y}_{it}=(y_{1it},\dots,y_{Kit})^{T}$ represent the vector of case counts for all diseases in area $i$ during the time interval (t1,t](t-1,t], where $y_{kit}$ represents the reported case counts for disease $k$. We will first present our model without zero inflation and then extend to the zero-inflated case in Section 2.1. Since we are mainly interested in comparing the diseases, we condition on the total number of cases and model $\bm{y}_{it}$ using a multinomial distribution,
@@
where $\text{total}_{it}=\sum_{k=1}^{K}y_{kit}$ and $\bm{y}_{t-1}=(\bm{y}_{1(t-1)},\dots,\bm{y}_{N(t-1)})^{T}$ is the vector of all case counts observed at the previous time point, giving the model an autoregressive structure. In (1), $\bm{\pi}_{it}=(\pi_{1it},\dots,\pi_{Kit})^{T}$ represents the expected relative distribution of the disease counts at time $t$ in area $i$ given the previously observed counts $\bm{y}_{t-1}$. If conditions in (t1,t](t-1,t] favor some diseases over others, then the share of the favored diseases should grow relative to those less favored. Therefore, we model the odds of a new case being from disease $k$ relative to the baseline disease (relative odds) as,
@@
for $k=2,\dots,K$, where we assume disease 1 is the baseline disease, and we add 1 to the top and bottom to avoid dividing by zero, adding such constants is common in autoregressive count models (Liboschik et al., 2017; Fritz and Kauermann, 2022). The parameters $\zeta_{k}$ and $\zeta_{1}$, assumed to be between 0 and 1, are meant to account for nonhomogeneous mixing (Wakefield et al., 2019). That is, as cases increase, there tends to be a dampening effect on transmission as individuals avoid infection and governments implement interventions. In (2), $\lambda_{kit}>0$, which we link to covariates below, is a parameter that measures the favorability of conditions for the transmission of disease $k$ relative to the baseline disease during (t1,t](t-1,t]. For instance, if $\lambda_{kit}>1$ ($\lambda_{kit}<1$) then we would expect the share of disease $k$ to grow (shrink) relative to the baseline disease in (t1,t](t-1,t], after adjusting for nonhomogeneous mixing with $\zeta_{k}$ and $\zeta_{1}$. If $\zeta_{k}\approx\zeta_{1}<<1$, like for our motivating example in Section 4, then this means that if $\lambda_{kit}>1$, the share of disease $k$ will grow initially from an equal share. However, as disease $k$ becomes more dominant, the share will be pulled back towards equality as individuals avoid infection from the dominant disease.
@@
$\log(\lambda_{kit})=\alpha_{0k}+\alpha_{0ki}+\bm{x}_{kit}^{T}\bm{\alpha}_{k}+\phi_{kit},$ (3)
where $\alpha_{0ki}\sim N(0,\sigma_{k}^{2})$ is a normal random intercept meant to account for between area differences and $\bm{x}_{kit}$ is a vector of covariates that might affect the favorability of disease $k$ relative to the baseline disease. We also considered a spatially correlated intrinsic conditional auto-regressive (ICAR) model for the intercepts, $\alpha_{0ki}\sim\text{ICAR}(\sigma_{k,\text{ICAR}}^{2})$ (Besag et al., 1991). However, for our motivating example in Section 4, while our proposed model converged, we were unable to obtain convergence for some of the comparison models when using ICAR intercepts. As the results of the normal random intercept and ICAR models were very similar, see SM Section 11, we decided to use normal random intercepts in Section 4 to ensure that all comparisons were fair. We give the results of the ICAR model applied to our motivating example in SM Section 11, and the code is available on GitHub https://github.com/Dirk-Douwes-Schultz/MS_ZIARMN_code. We note that the number of cases in neighboring areas may affect the favorability of the diseases, as individuals will mix with those in neighboring areas (Grenfell et al., 2001). Therefore, following Bauer and Wakefield (2018) and Bracher and Held (2022), we allow functions of disease cases in neighboring areas, such as the prevalence across neighboring areas, to be added to the covariate vector $\bm{x}_{kit}$ to account for the geographical spread of the disease. Also, we allow previous counts of other diseases, e.g., $\log(y_{ji(t-1)}+1)$ for $j\neq k$, to be added to $\bm{x}_{kit}$ to account for potential disease interactions (Freitas et al., 2019), see Section 4 for an example.
We model the random effects $\phi_{kit}$ using a multivariate normal distribution (Xia et al., 2013; Zeng et al., 2022),
@@
where $\bm{\Sigma}$ is a K1K-1 by K1K-1 variance-covariance matrix. These random effects are mainly meant to account for overdispersion, which is a very common issue when modeling multinomial counts (Xia et al., 2013). Since $\phi_{kit}$ is correlated across diseases, they also allow for a more flexible correlation structure between the individual disease counts compared to the standard multinomial distribution; see SM Section 2.
@@
There is an epidemiological interpretation of the parameters in Equations (1)-(4). As we show in SM Section 1, if all diseases follow a Reed-Frost chain binomial SIR model (Abbey, 1952; Vynnycky, 2010; Bauer and Wakefield, 2018) with roughly the same serial interval, the time between successive generations of cases, then we can derive Equations (1)-(4). Under this derivation, $\lambda_{kit}$ in (2) represents the ratio of the effective reproduction number of disease $k$ and the baseline disease. The effective reproduction number is an important measure of disease transmission in epidemiology, representing the average number of new infections produced by a single infectious individual in the current population before they recover (Vynnycky, 2010). The covariate effect $\alpha_{kj}$ then represents the difference between the effect of covariate $x_{kitj}$ on the effective reproduction number of disease k and the baseline disease; see SM Section 1 for more details.
@@
While the Reed-Frost model is commonly used (Wakefield et al., 2019), it makes assumptions that are not appropriate for many diseases (Abbey, 1952), and the model can be sensitive to underreporting and reporting delay when it comes to estimating the effective reproduction number (Bracher and Held, 2021; Quick et al., 2021). Regardless, the Reed-Frost derivation does reveal an important source of potential confounding due to differences in the susceptible populations, i.e., the population of individuals who are not immune to the disease. From the Reed-Frost derivation, see SM Equation (8), ideally, we would add $\log(\delta_{ki(t-1)}/\delta_{1i(t-1)})$ to $\log(\lambda_{kit})$ in Equation (3), where $\delta_{ki(t-1)}$ is the size of the susceptible population for disease $k$ in area $i$ at time t1t-1. This makes sense intuitively; if a disease has a larger susceptible population compared to another, then transmission for that disease will naturally be favored. The effect of any covariates correlated with the ratio of the susceptible populations could then be confounded. We conduct a sensitivity analysis to attempt to control for this for our motivating example in SM Section 9.
We will call the model defined by (1)-(4) the autoregressive multinomial (ARMN) model. As detailed in the introduction, the ARMN model is not appropriate if some of the diseases are absent in an area, as it will always assign a positive probability to cases being reported. In a simulation study in SM Section 4, we show that ignoring long periods of disease absence leads to poor coverage of the 95% credible intervals and posteriors centered far from the actual parameter values. In the next subsection, we will keep the definitions from above and extend the ARMN model to deal with long periods of disease absence by adapting the proposal of Zeng et al. (2022) to our space-time setting.